A grocer plans to mix candy that sells for $2 per pound with candy
that sells for $4 per pound to get a mixture that sells for $2.75 per pound. How much of each type of candy should she use if she wants 40 pounds of the mixture?

Answer by lwsshak3(11628)

A grocer plans to mix candy that sells for $2 per pound with candy
that sells for $4 per pound to get a mixture that sells for $2.75 per pound. How much of each type of candy should she use if she wants 40 pounds of the mixture?
***
let x=amount of $2 candy to use
40-x=amount of $4 candy to use
..
2x+4(40-x)=40(2.75)
2x+160-4x=110
2x=50
x=25
40-x=15
amount of $2 candy to use=25 lbs
amount of $4 candy to use=15 lbs